# Inferring – Evolution from the Gravitational Wave Background

### v1

@@ 1 Introduction
In the local universe, it has been shown that there is a strong correlation between host galaxy bulge mass and central SMBH mass, making it possible to accurately predict SMBH masses in nearby galaxies. Direct, independent SMBH mass measurements are difficult outside the local universe, however, and so very few constraints exist for this relation for . High redshift SMBH mass measurements are hindered by telescope resolution limits meaning that masses must be inferred indirectly. Because of their extreme luminosity, AGN can be used to place useful constraints on SMBH masses at redshifts as high as (J. Jeon et al., 2025; Á. Bogdán et al., 2024; R. Maiolino et al., 2024a; L. Napolitano et al., 2024). From the AGN luminosity function, one can infer the BHMF, but assumptions about, e.g., radiative efficiency and AGN fraction lead towards large uncertainties (X. Shen et al., 2020). Furthermore, these types of surveys are frequently magnitude limited and are thus subject to observational biases (T. R. Lauer et al., 2007).
@@
While AGN studies inform SMBH populations, modeling the GWB calls for a simpler BHMF framework. The observed correlation between galaxy bulge mass and SMBH mass is known as the – relation (J. Kormendy & L. C. Ho, 2013; N. J. McConnell & C.-P. Ma, 2013). To approximate SMBH number density at higher redshifts, one can convolve the local – relation with the galaxy stellar mass function (GSMF), which is well-observed for the redshifts where most of the GWB signal is expected to be produced ( A. Sesana et al., 2004; G. Agazie et al., 2023b; J. Leja et al., 2020). This method assumes the – relation to be unchanging with time, an assumption with ambiguous observational support. G. Agazie et al. (2023b) derived physical quantities from fitting the NANOGrav 15-year PTA data and their results suggested that the data are best described with a number density of SMBHs, with , that is notably greater than predictions from current models (such as J. Kormendy & L. C. Ho, 2013; N. J. McConnell & C.-P. Ma, 2013). Since then, other studies have used a variety of techniques finding similar results (e.g., G. Sato-Polito et al., 2024; Y. Chen et al., 2024; E. R. Liepold & C.-P. Ma, 2024). Additionally, there exist other scaling relations, such as the – relation, which connects SMBH mass to galaxy velocity dispersion (L. Ferrarese & D. Merritt, 2000; K. Gebhardt et al., 2000). It has been shown that the – and – relations make different predictions for the BHMF at high redshift such that the predicted GWB is higher when using – (C. Matt et al., 2023; J. Simon, 2023). This difference in BHMF predictions may be due to evolution in one or both relations that is not accounted for. Via the three-way relationship between galaxy stellar mass, radius, and velocity dispersion (A. de Graaff et al., 2021), the – relation naturally accounts for galaxy downsizing (A. van der Wel et al., 2014) and may therefore be a more fundamental probe of the SMBH’s gravitational potential well than – (R. C. E. van den Bosch, 2016; J. H. Cohn et al., 2025). This combination of findings suggests that methods of predicting SMBH number density are in need of revision.
@@ 2.1 Semi-Analytic Modeling
We use holodeck (details can be found in section 3 of G. Agazie et al., 2023b)111Also see https://github.com/nanograv/holodeck. to synthesize populations of massive black holes. We start by convolving the galaxy stellar mass function from J. Leja et al. (2020) with the SMBH–galaxy mass scaling relation from (J. Kormendy & L. C. Ho, 2013). The specifics of this implementation are detailed in sections 2.2 and 2.3. We then apply galaxy merger rates, and SMBH binary hardening models to solve for the number density of SMBH binaries emitting gravitational waves in frequencies detectable by PTAs, i.e., the GWB spectrum predicted from the input parameters. Previously, G. Agazie et al. (2023b) calculated galaxy merger rates from galaxy pair fractions and merger times following the prescription due to (S. Chen et al., 2019a). For this work, we instead use the galaxy merger rate prescription due to V. Rodriguez-Gomez et al. (2015). A detailed description and graphic of the holodeck workflow can be found in section 3 of G. Agazie et al. (2023b).
To determine our best-fit parameters for each model, we first sampled 20,000 times from a set of astrophysically-motivated prior distributions with a Latin hypercube. We generated and averaged over 100 realizations for each set of sampled parameters to account for the Poisson sampling in the GWB spectrum calculations. Our sample size is higher than was used in G. Agazie et al. (2023b), who used Gaussian process interpolation for their posterior parameter estimation. Because of our larger parameter space (up to 30 parameters versus their 6), it is not computationally feasible to follow their same methods (N. Laal et al., 2025). Instead, our finer sampling makes it so that we can compare each model to the spectrum without the need for Gaussian process interpolation between the models. We compare our models to the 15 year HD-correlated free-spectrum representation of the GWB (specifically, we use the “HD-w/MP+DP+CURN” case which is detailed further in G. Agazie et al., 2023a; W. G. Lamb et al., 2023). For each GWB frequency, we evaluate the goodness of fit for a given modeled spectrum by comparing the value of each model to the probability density distribution of the data to determine how well each model fits the data. This process produces results with the same fidelity as G. Agazie et al. (2023b) with a higher computational efficiency (see their Appendix C, also our figures in Appendix C).
@@ 2.3 The Galaxy Stellar Mass Function
G. Agazie et al. (2023b) used the single-Schechter (P. Schechter, 1976) GSMF prescription due to S. Chen et al. (2019b). For this work, we predominantly use the GSMF due to J. Leja et al. (2020), which is defined as a double-Schechter function offering a more accurate description of the mass function of the total galaxy population (e.g., D. J. McLeod et al., 2021, and references therein). This GSMF is strongly supported by observational data and has an explicitly defined evolution making it possible to reliably calculate the GSMF at any specific redshift, even extrapolating outside the range of the data (). J. Leja et al. (2020) find that the redshift evolution of the characteristic mass and each of the density normalization terms is best described by a Gaussian. Following their methods (see their equations 14 and B 1-4), we define the full functional form of the GSMF to be
@@ 2.4 The Models
The largest models include 29 parameters (30 with ). Of these parameters, the GSMFs make up 11 variables, five that describe the local GSMF and six that contain information about the GSMF evolution. To help understand the degeneracy between a changing GSMF versus – relation we consider three different cases of sampling the GSMF: one in which we sample all 11 variables (Le11ne and Le11ev), one in which we only sample three local values (Le03ne and Le03ev), and one with all GSMF parameters fixed to their fiducial values (Le00ne Le00ev). The three GSMF values that are varied in Le03ne and Le03ev are , , and . For all but 2 of our 8 models we use the Gaussian posteriors from J. Leja et al. (2020) as priors for our GSMF and we assume that these parameters are independent of each other. For the last two models, we use our explicitly evolving version of the E. R. Liepold & C.-P. Ma (2024) GSMF, which we described previously in section 2.3 and also in appendix section A.
Each model we present is given a short-hand name to indicate the important features. Each name is 6 total characters to convey three pieces of information: The first two characters indicate which version of the GSMF is being used where Le is J. Leja et al. (2020) and LM is E. R. Liepold & C.-P. Ma (2024). The middle two characters indicate how many of the GSMF parameters (out of 11) were sampled where the number of parameters fixed to their fiducial values is 11 minus that number. The final two characters indicate whether the – amplitude was allowed to evolve (ev) or not (ne). This information is summarized in Table 1.
@@
Because the models with evolving – relations are a strict parametric super set of those without (i.e., the models are nested), we performed Savage-Dickey ratio tests to quantify the significance of the evolving model over the fixed model (J. M. Dickey, 1971; E.-J. Wagenmakers et al., 2010). For all our figures and calculations we present the results from fits to the first 5 frequency bins. Our conclusions are not sensitive to the number of frequency bins we fit to.
@@ 3 Results
A summary of the results for the posterior values for are shown in Table 2 and Figure 1. The best-fit spectra for all our models are presented in Figure 3. We find that, in general, a large number of massive SMBHs are needed to reproduce the GWB. When is allowed to vary, the majority of the posterior distribution is positive with a median value of across all models and for models with statistically significant evolution. We additionally find that, for models with fixed to be 0, this high number density of massive SMBHs can be modeled either through a top-heavy GSMF (between 0.2 to 3 dex higher in number density for and compared to the observed GSMF, see Figures 4 and 5) or a high – amplitude (with typically compared to the locally observed value of 8.69, J. Kormendy & L. C. Ho, 2013). We first present our full-sample model containing posterior distributions and best-fit spectra for all 29 (30 with ) parameters. Then, to highlight the effects of the degeneracy between parameters, we repeat this process with a subset of sampled parameters.
@@ 3.1.1 Le11ne and Le11ev
For the majority of parameters, the posterior distributions recover the priors for both Le11ne and Le11ev. In the top row of Figure 2 we highlight 8 parameters from these models; complete corner plots can be found in Appendix B. The non-evolving model, Le11ne shows slight deviations towards higher values for three of the the GSMF characteristic mass and normalization evolutionary parameters (, , and ). The two posterior distributions with the largest deviations from the priors are both of the evolutionary parameters for the characteristic stellar mass. When the – relation is allowed to evolve, the posterior distributions for Le11ev are more consistent with the priors for all GSMF parameters, though there is still a mild offset in the same direction as Le11ne for . The posterior distribution for is 87.2% positive with a median value of 0.85. A Savage-Dickey ratio test in favor of the evolving model returns a value of 1.3, which indicates that this fit is consistent with no – evolution.
The local GSMF values are recovered in both models, but the non-evolving model posteriors return generally larger values for the evolutionary parameters. Notably, 93% of the posterior distribution for is greater than the prior median and 47% of the distribution falls above the prior upper limit. Higher values for and produce greater number densities of galaxies at higher redshifts, while maintaining consistency with the fiducial local number density. In particular, larger values for the characteristic mass produce more massive galaxies and therefore a greater number of massive SMBHs.
@@ 3.2 Fixed GSMF Evolution Models: Le03ne and Le03ev
The first of the sub-models that we discuss are Le03ne and Le03ev. Each of these is identical to Le11ne and Le11ev respectively except the six evolutionary parameters (, , , , , and ) alongside and for the GSMF are fixed to the posterior values given by J. Leja et al. (2020). To focus on the degeneracy between GSMF and – parameters we additionally fixed all parameters for galaxy merger rates and bulge fractions to their fiducial values. These models, therefore, sample a total of 8 parameters (9 with ).
@@
When the – relation is allowed to evolve, the posterior distributions for all GSMF parameters are in good agreement with the priors, even more than with Le11ev. Interestingly, the distribution for , for this model, is 99.3% positive with a Savage-Dickey ratio of 24.6. The standard deviation of this distribution is also significantly reduced indicating a higher confidence of positive evolution.
Nearly all samples from the model that did not include – evolution tend towards lower GWB amplitudes. In the second panel of Figure 3 we see that the shape of the GWB spectrum is nearly straight with the best fit model lying below the bulk of the data for the majority of the frequency bins. The Le03ev model, on the other hand, while lower in GWB amplitude than Le11ev at the high frequency end, is still consistent with the PTA data.
@@ 3.3 Fixed GSMF Models: Le00ne and Le00ev
The GWB spectra share similar characteristics to the Le03ne and Le03ev models with the non-– evolving models consistently lying below the majority of the PTA data and the evolving – models falling closer to the center of the violins in more bins, but lying slightly above the median. We note that the best-fit spectrum for Le03ne has the lowest likelihood of all 8 spectra, an effect we detail further in Appendix C.

### v2

@@ 1 Introduction
In the local universe, it has been shown that there is a strong correlation between host galaxy bulge mass and central SMBH mass, making it possible to accurately predict SMBH masses in nearby galaxies (J. Kormendy & L. C. Ho, 2013; N. J. McConnell & C.-P. Ma, 2013). Direct, independent SMBH mass measurements are difficult outside the local universe, however, and so very few constraints exist for this relation for . High redshift SMBH mass measurements are hindered by telescope resolution limits meaning that masses must be inferred indirectly. Because of their extreme luminosity, AGN can be used to place useful constraints on SMBH masses at redshifts as high as (J. Jeon et al., 2025; Á. Bogdán et al., 2024; R. Maiolino et al., 2024a; L. Napolitano et al., 2024). From the AGN luminosity function, one can infer the BHMF, but assumptions about, e.g., radiative efficiency and AGN fraction lead towards large uncertainties (X. Shen et al., 2020). Furthermore, these types of surveys are frequently magnitude limited and are thus subject to observational biases (T. R. Lauer et al., 2007).
@@
While AGN studies inform SMBH populations, modeling the GWB calls for a simpler BHMF framework. The observed correlation between galaxy bulge mass and SMBH mass is known as the – relation (J. Kormendy & L. C. Ho, 2013; N. J. McConnell & C.-P. Ma, 2013). To approximate SMBH number density at higher redshifts, one can convolve the local – relation with the galaxy stellar mass function (GSMF), which is well-observed for the redshifts where most of the GWB signal is expected to be produced ( A. Sesana et al., 2004; G. Agazie et al., 2023b; J. Leja et al., 2020). This method assumes the – relation to be unchanging with time, an assumption with ambiguous observational support. It expected that analyses of the GWB can be used to place constraints on the underlying population of galaxies and black holes (e.g., J. Simon & S. Burke-Spolaor, 2016; D. Izquierdo-Villalba et al., 2022; S. Bonoli et al., 2025). G. Agazie et al. (2023b) derived physical quantities from fitting the NANOGrav 15-year PTA data and their results suggested that the data are best described with a number density of SMBHs, with , that is notably greater than predictions from current models (such as J. Kormendy & L. C. Ho, 2013; N. J. McConnell & C.-P. Ma, 2013). Since then, other studies have used a variety of techniques finding similar results (e.g., G. Sato-Polito et al., 2024; Y. Chen et al., 2024; E. R. Liepold & C.-P. Ma, 2024). Additionally, there exist other scaling relations, such as the – relation, which connects SMBH mass to galaxy velocity dispersion (L. Ferrarese & D. Merritt, 2000; K. Gebhardt et al., 2000). It has been shown that the – and – relations make different predictions for the BHMF at high redshift such that the predicted GWB is higher when using – (C. Matt et al., 2023; J. Simon, 2023). This difference in BHMF predictions may be due to evolution in one or both relations that is not accounted for. Via the three-way relationship between galaxy stellar mass, radius, and velocity dispersion (A. de Graaff et al., 2021), the – relation naturally accounts for galaxy downsizing (A. van der Wel et al., 2014) and may therefore be a more fundamental probe of the SMBH’s gravitational potential well than – (R. C. E. van den Bosch, 2016; J. H. Cohn et al., 2025). This combination of findings suggests that methods of predicting SMBH number density are in need of revision.
@@ 2.1 Semi-Analytic Modeling
We use holodeck (details can be found in section 3 of G. Agazie et al., 2023b)111Also see https://github.com/nanograv/holodeck. to synthesize populations of massive black holes. We start by convolving the galaxy stellar mass function from J. Leja et al. (2020) with the SMBH–galaxy mass scaling relation from (J. Kormendy & L. C. Ho, 2013). The specifics of this implementation are detailed in sections 2.2 and 2.3. We then apply galaxy merger rates, and SMBH binary hardening models to solve for the number density of SMBH binaries emitting gravitational waves in frequencies detectable by PTAs, i.e., the GWB spectrum predicted from the input parameters. Previously, G. Agazie et al. (2023b) calculated galaxy merger rates from galaxy pair fractions and merger times following the prescription due to (S. Chen et al., 2019a). For this work, we instead use the galaxy merger rate prescription due to V. Rodriguez-Gomez et al. (2015), which provides a more astrophysically motivated model for mergers. This change does not significantly affect the shape or amplitude of the GWB spectrum (see Figure 10 in Appendix A). A detailed description and graphic of the holodeck workflow can be found in section 3 of G. Agazie et al. (2023b).
To determine our best-fit parameters for each model, we first sampled 20,000 times from a set of astrophysically-motivated prior distributions with a Latin hypercube. We generated and averaged over 100 realizations for each set of sampled parameters to account for the Poisson sampling in the GWB spectrum calculations. Our sample size is higher than was used in G. Agazie et al. (2023b), who used Gaussian process interpolation for their posterior parameter estimation. Because of our larger parameter space (up to 30 parameters versus their 6), it is not computationally feasible to follow their same methods (N. Laal et al., 2025). Instead, our finer sampling makes it so that we can compare each model to the spectrum without the need for Gaussian process interpolation between the models. We compare our models to the 15 year Hellings and Downs (HD)-correlated free-spectrum representation of the GWB (W. G. Lamb et al., 2023). Specifically, we use the “HD-w/MP+DP+CURN” model which included both monopole (MP) and dipole (DP) correlated red noise in addition to the common uncorrelated red noise (CURN) (see additional details in G. Agazie et al., 2023b). For each GWB frequency, we evaluate the goodness of fit for a given modeled spectrum by comparing the value of each model to the probability density distribution of the data to determine how well each model fits the data. This process produces results with the same fidelity as G. Agazie et al. (2023b) with a higher computational efficiency (see their Appendix C, also our figures in Appendix C).
Throughout this paper we refer to the “likelihood” of the models which we calculate in our fitting process. These likelihoods represent how well models fit the GWB data, but do not contain information on how well the models agree with observational constraints from EM datasets. Therefore we additionally quantify goodness of fit to, e.g. the GSMF, with statistical tests such as the Kullback-Leilbler divergence (, S. Kullback & R. A. Leibler, 1951, see Section 3), and a quantity, , which we define in Section 3.4 to evaluate the agreement between our models and the observed GSMF.
@@ 2.3 The Galaxy Stellar Mass Function
G. Agazie et al. (2023b) used the single-Schechter (P. Schechter, 1976) GSMF prescription due to S. Chen et al. (2019b). For this work, we predominantly use the GSMF due to J. Leja et al. (2020), which is defined as a double-Schechter function offering a more accurate description of the mass function of the total galaxy population (e.g., D. J. McLeod et al., 2021, and references therein). This GSMF is strongly supported by observational data and has an explicitly defined evolution making it possible to reliably calculate the GSMF at any specific redshift. This model is reliable within the range of the data () though we extrapolate the model to higher redshifts in holodeck for defining our populations. The reliability of extrapolating this model is uncertain, but the majority of the GWB signal is dominated by lower redshifts (G. Agazie et al., 2023b; A. Sesana et al., 2004) so we do not expect our results to be influenced by any error that may be introduced by this extrapolation. J. Leja et al. (2020) find that the redshift evolution of the characteristic mass and each of the density normalization terms is best described by a Gaussian. Following their methods (see their equations 14 and B 1-4), we define the full functional form of the GSMF to be
@@ 2.4 The Models
The largest models include 29 parameters (30 with ). Of these parameters, the GSMFs make up 11 variables, five that describe the local GSMF and six that contain information about the GSMF evolution. For six of our eightmodels the priors for sampled GSMF parameters are based on the posterior fit values from J. Leja et al. (2020). Our fiducial parameters are therefore the median value of these distributions. For two of our eightwe use the same functional form of the GSMF, but with local values due to E. R. Liepold & C.-P. Ma (2024) and evolutionary parameters we define for this work (see appendix A). To help understand the degeneracy between a changing GSMF versus – relation we consider three different cases of sampling the J. Leja et al. (2020) GSMF: one in which we sample all 11 variables (Le11ne and Le11ev), one in which we only sample three local values (Le03ne and Le03ev), and one with all GSMF parameters fixed to their fiducial values (Le00ne Le00ev). The three GSMF values that are varied in Le03ne and Le03ev are , , and . For all but two of our 8 models we use the Gaussian posteriors from J. Leja et al. (2020) as priors for our GSMF and we assume that these parameters are independent of each other. For the last two models, we use our explicitly evolving version of the E. R. Liepold & C.-P. Ma (2024) GSMF, which we described previously in section 2.3 and also in appendix section A.
Each model we present is given a short-hand name to indicate the important features. Each name is six total characters to convey three pieces of information: The first two characters indicate which version of the GSMF is being used where Le is J. Leja et al. (2020) and LM is E. R. Liepold & C.-P. Ma (2024). The middle two characters indicate how many of the GSMF parameters (out of 11) were sampled where the number of parameters fixed to their fiducial values is 11 minus that number. The final two characters indicate whether the – amplitude was allowed to evolve (ev) or not (ne). This information is summarized in Table 1.
@@
Because the models with evolving – relations are a strict parametric super set of those without (i.e., the models are nested), and since our prior for is uninformative, we performed Savage-Dickey ratio tests to quantify the significance of the evolving model over the fixed model (J. M. Dickey, 1971; E.-J. Wagenmakers et al., 2010). For all our figures and calculations we present the results from fits to the first 5 frequency bins. Our conclusions are not sensitive to the number of frequency bins we fit to.
@@ 3 Results
A summary of the results for the posterior values for are shown in Table 2 and Figure 1. In Figure 2 we highlight parameters of interest for which we calculate Kullback-Leibler divergences, , between posterior and prior distributions (S. Kullback & R. A. Leibler, 1951). The best-fit spectra for all our models are presented in Figure 3. We find that, in general, a large number of massive SMBHs are needed to reproduce the GWB. When is allowed to vary, the majority of the posterior distribution is positive with a median value of across all models and for models with statistically significant evolution. We additionally find that, for models with fixed to be 0, this high number density of massive SMBHs can be modeled either through a top-heavy GSMF (between 0.2 to 3 dex higher in number density for and compared to the observed GSMF, see Figures 4 and 5) or a high – amplitude (with typically compared to the locally observed value of 8.69, J. Kormendy & L. C. Ho, 2013). We first present our full-sample model containing posterior distributions and best-fit spectra for all 29 (30 with ) parameters. Then, to highlight the effects of the degeneracy between parameters, we repeat this process with a subset of sampled parameters.
@@ 3.1.1 Le11ne and Le11ev
For the majority of parameters, the posterior distributions recover the priors for both Le11ne and Le11ev. In the top row of Figure 2 we highlight eight parameters from these models; we discuss the effects of the remaining parameters in appendix C and complete corner plots can be found in Appendix D. The non-evolving model, Le11ne shows slight deviations towards higher values for three of the the GSMF characteristic mass and normalization evolutionary parameters (, , and ). The two posterior distributions with the largest deviations from the priors are both of the evolutionary parameters for the characteristic stellar mass. When the – relation is allowed to evolve, the posterior distributions for Le11ev are more consistent with the priors for all GSMF parameters, though there is still a mild offset in the same direction as Le11ne for . The posterior distribution for is 87.2% positive with a median value of 0.85. A Savage-Dickey ratio test in favor of the evolving model returns a value of 1.3, which indicates that this fit is consistent with no – evolution.
The local GSMF values are recovered in both models, but the non-evolving model posteriors return generally larger values for the evolutionary parameters. Higher values for and produce greater number densities of galaxies at higher redshifts, while maintaining consistency with the fiducial local number density. In particular, larger values for the characteristic mass produce more massive galaxies and therefore a greater number of massive SMBHs.
We use to quantify the deviation of the posteriors from the priors for Le11ne and Le11ev. Two distributions with are equivalent while indicate disagreement. For all the parameters shown in Figure 2, the value of for Le11ne is equivalent or greater than for Le11ev. This means that the posterior distributions for Le11ne diverge from the priors more than in Le11ev suggesting that the posterior distributions for Le11ev are in better agreement with the observed GSMF and local – relation.
@@ 3.2 Fixed GSMF Evolution Models: Le03ne and Le03ev
The first of the sub-models that we discuss are Le03ne and Le03ev. Each of these is identical to Le11ne and Le11ev respectively except the six evolutionary parameters (, , , , , and ) alongside and for the GSMF are fixed to the posterior values given by J. Leja et al. (2020). To focus on the degeneracy between GSMF and – parameters we additionally fixed all parameters for galaxy merger rates and bulge fractions to their fiducial values. These models, therefore, sample a total of eight parameters (nine with ).
@@
When the – relation is allowed to evolve, the posterior distributions for all GSMF parameters are in good agreement with the priors. Interestingly, the distribution for , for this model, is 99.3% positive with a Savage-Dickey ratio of 24.6. The standard deviation of this distribution is also significantly reduced indicating a higher confidence of positive evolution.
Similarly to Le11ne versus Le11ev, the between the priors and posteriors for Le03ne is equivalent or greater than for Le03ev. As mentioned earlier, the posterior distributions for and diverge from the priors more significantly in Le03ne than in Le11ne. In Le11ne the two distributions with the highest divergence were , and . These values are fixed in Le03ne and so this model compensates by increasing , and . We do not see a notable change to the values for these distributions in Le00ev versus Le11ev.
Nearly all samples from the model that did not include – evolution tend towards lower GWB amplitudes. In the second panel of Figure 3 we see that the shape of the GWB spectrum is nearly straight. Except for the two right-most frequency bins, the best-fit spectrum is equivalent to the upper limit on the models meaning that no models produced a GWB amplitude higher than the best-fit model for the higher-frequencies. The best-fit model here is roughly consistent (though on the low side) with four of the five frequency bins we used for fitting. The Le03ev model, on the other hand, while lower in GWB amplitude than Le11ev at the high frequency end, is still consistent with the PTA data.
@@ 3.3 Fixed GSMF Models: Le00ne and Le00ev
The GWB spectra share similar characteristics to the Le03ne and Le03ev models with the non-– evolving models consistently lying below the majority of the PTA data and the evolving – models falling closer to the center of the violins in more bins, but lying slightly above the median. We note that the best-fit spectrum for Le00ne has the lowest likelihood of all 8 spectra indicating it is the worst fit to the GWB, an effect we detail further in Appendix C.